OpenStudy (anonymous):

the derivative of square root of t is 2t?

OpenStudy (anonymous):

no it's 1/2t

OpenStudy (anonymous):

because square root is the same as t^1/2 and you bring that to the front in derivatives.

OpenStudy (anonymous):

\[\frac{1}{2\sqrt{t}}\]

OpenStudy (anonymous):

oh yeah hehe brain fart

OpenStudy (anonymous):

can anyone help me in product rule?

OpenStudy (anonymous):

on*

OpenStudy (anonymous):

Yeah i can.

OpenStudy (anonymous):

okay so I have \[2x ^{3}e ^{7x}\] and i have to use product rule. So almost at the final stage I got (2x^3)(7e^(7x))+(e^7x)(6x^2)...but I'm a bit confused on how to multiply them and simplify, especially witth the exponents

OpenStudy (anonymous):

Ok i will start from the beginning so i dont get confused lol.

OpenStudy (anonymous):

\[(6x^2)(e^{7x}) + (2x^{3})(e^{7x})\]

OpenStudy (anonymous):

yes thats what i got, now how to simplify

OpenStudy (anonymous):

I might of taken the derivative of \[e^{7x}\] wrong.. let me check before i move on.

OpenStudy (anonymous):

Yeah i did. Let me correct it.

OpenStudy (anonymous):

oh yeah u did, it was suppose to be 7 in front

OpenStudy (anonymous):

\[(6x^{2})(e^{7x}) + (3x^{2})(7e^{7x})\]

OpenStudy (anonymous):

Ok we can factor out a e^7x

OpenStudy (anonymous):

okay wait, its suppose to be uv'+vu' form...

OpenStudy (anonymous):

\[e^{7x}[6x^{2} + 3x^{2}*7]\]

OpenStudy (anonymous):

The order doesnt matter in the product rule.

OpenStudy (anonymous):

okay but shouldnt the first parentheses ont he second set be \[2x ^{?}\] not 3x^2

OpenStudy (anonymous):

i mean 3x^3

OpenStudy (anonymous):

ahh i bumped the three and i ment the 2.. Lets see if i can get it right this time >.< lol

OpenStudy (anonymous):

\[(6x^2)(e^{7x})+(2x^3)(7e^{7x})\]

OpenStudy (anonymous):

Ok did i get it right this time?? (i hope lol )

OpenStudy (anonymous):

Factor out e^7x \[e^{7x}[6x^2 + 14x^3]\] and you can also factor out a 2x^2 \[e^{7x}2x^2[3 + 7x]\] And that is the answer!

OpenStudy (anonymous):

http://www.wolframalpha.com/input/?i=%282x%5E3%29%28e%5E%287x%29%29&t=mfftb01 This confirms i didnt screw it up this time lol

OpenStudy (anonymous):

okay thanks! can you help me with one more?

OpenStudy (anonymous):

Sure

OpenStudy (anonymous):

f(x)=\[-x ^{2}e ^{5\sin(x)}\]

OpenStudy (anonymous):

Ok \[(-2x)(e^{5\sin(x)}) + (-x^2)(5e^{5\sin(x)}\cos(x))\]

OpenStudy (anonymous):

Factor out -xe^5sin(x)\[-e^{5\sin(x)}x[2 + 5xcos(x)]\]

OpenStudy (anonymous):

okay for the second set, it couldnt have been written like 3cos(x)e^(5sinx)?

OpenStudy (anonymous):

or does the cos have to be beside 5e adn its exponent

OpenStudy (anonymous):

by 3 i meatnt 5

OpenStudy (anonymous):

it would be \[\cos(x)5e^{5\sin(x)}\] if you wanted the cosine in the front. The way i have it written is correct though and its just force of habbit when i use the chain rule i write the derivative of the inner function last.

OpenStudy (anonymous):

okay, well we havent been taught the chain rule yet. we are next week.

OpenStudy (anonymous):

Wow. You need the chain rule to solve the last one lol.. Are you working ahead in class or something or is your teacher that mean lol :P

OpenStudy (anonymous):

perhaps she didnt realize it required the chain rule. I don't know. She's not mean, but she expects us to know stuff :O ...but y exactly do we need the chain rule for it?

OpenStudy (anonymous):

Because the e^5sin(x) in a composite function. And to find the derivative of a composite function you need the Chain Rule.

OpenStudy (anonymous):

Its pretty simple to learn, i can teach it to you real quick if you want.

OpenStudy (anonymous):

oh actually we skipped to chapter 5 for those involiving e....natural exponential functions and how to find their derivatives....(is that the same or different?)

OpenStudy (anonymous):

Same or different as what??

OpenStudy (anonymous):

the same as finding the derivative with the chain rule.. because this is the form to gett he derivatives involving e f'(u)=(e^u)du

OpenStudy (anonymous):

Yeah its kinda like the chain rule.

OpenStudy (anonymous):

okay. so what is the chain rule? Can you teach me? I'll gladly appreciate it. That way i can be ahead in knowing.

OpenStudy (anonymous):

Ok.. Chain rule is used to find the derivative of Composite Functions in the form \[f(g(x))\] ( f composed with g(x) ).. Where f(x) is the outer function and g(x) is the inner function.

OpenStudy (anonymous):

So say we have the function \[y = (2x + 1)^2\]

OpenStudy (anonymous):

The outer function is \[f(x) = x^2\] and the inner function is \[g(x) = 2x + 1\]. F is composed with g(x). Do you understand so far?

OpenStudy (anonymous):

yes I do but why is the outer function xsquared?

OpenStudy (anonymous):

Its like doing the composition of a function in reverse. You have \[f(g(x)) = (2x + 1)^2\] What is f(x) and what is g(x).. f(x) is the ^2 and g(x) is 2x + 1

OpenStudy (anonymous):

ok got it

OpenStudy (anonymous):

i just use x because it is the most common variable used. It can be y^2 or z^2 ect...

OpenStudy (anonymous):

Ok.. before we move on, just to make sure you understand \[f(g(x)) = \sqrt{2x^2 + 3}\] What is f(x) (outer function) and what is g(x) (inner function)??

OpenStudy (anonymous):

inner function is 2x^2+3 and outer function is x^1/2?

OpenStudy (anonymous):

Good!

OpenStudy (anonymous):

Ok, if we have a composite function in the form f(g(x)) then to find the derivative you do the following. \[y \prime = f \prime(g(x))* g \prime (x)\]

OpenStudy (anonymous):

You take the derivative of the outer function and leave the inner function alone then multiply it by the derivative of the inner function.

OpenStudy (anonymous):

So going back to the example problem \[(2x+1)^2\] The outer function is \[f(x) = x^2\] and the inner function is \[g(x) = 2x + 1\]

OpenStudy (anonymous):

The chain rule first says to find the derivative of the outer function. So the derivative of \[x^2\] is \[2x\] and to put g(x) where x is so it will look like \[2(2x+1)\] then multiply it by the derivative of 2x + 1 which is 2. So it will look like this \[2(2x+1) * 2\]

OpenStudy (anonymous):

Make sense??

OpenStudy (anonymous):

yes yes

OpenStudy (anonymous):

Then after you simplify it you will end up with \[8x + 4\]

OpenStudy (anonymous):

Its pretty simple once you know how to do it :D

OpenStudy (anonymous):

hold on, i believe a question is coming in..

OpenStudy (anonymous):

okay got it! This is kind of awkward to ask but do you have a facebook page?

OpenStudy (anonymous):

You sure you understand? And i do.

OpenStudy (anonymous):

Yeah I understand completely what you told me. Do you think you can add me as a friend?

OpenStudy (anonymous):

Sure.

OpenStudy (anonymous):

okay my email is cald.dany@yahoo.com

OpenStudy (anonymous):

Do you have pickachu as a profile pic??

OpenStudy (anonymous):

hah yeah

OpenStudy (anonymous):

thats me